100 students are polled. 57 like spaghetti. 90 like pizza. 7 like neither. How many students like pizza and spaghetti.
Answer by checkley71(8403)

100-7=93 like pizza or spaghetti
90 like pizza & 57 like spaghetti
90+57=147 but we only have 93 in total that like pizza or spaghetti therefore:
147-93=54 that like pizza & spaghetti.
proof
---p-------s----ps-n code(p=pizza,s=spaghetti,ps=both pizza&spaghetti,n=none
(90-54)+(57-54)+54+7=100
36+3+54+7=100
100=100